OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x – y = 2 6x – 2y = 4

OpenStudy (anonymous):

This method requires that you rewrite one of the equations as y = ?? or x = ??. from there you substitute that equation for the variable into the 2nd equation. In this case , rewrite the first one and solve for Y. Once I see that, I will further explain next steps

OpenStudy (zpupster):

y=3x-2 6x-2y =4 6x-2(3x-2) = 4 6x-6x +4 = 4 4=4 this is a true statement

OpenStudy (anonymous):

zpupster - that is not the directions. I think what you were trying to say is that there is no solution. My question - is this consistent or inconsistent.

OpenStudy (anonymous):

Inconsistent?